If g:(2,0) = (2,00) is defined by g(x) Show that the function g is one-to- X-2 one and onto.
Answer #1

That is g(x_{1}) \neq g(x_{2}) . Therefore g is injective.

Now To prove onto. is onto. Suppore then for for any E (2,00), there existe ne (20) such that y=80) ntl y = 1+ 3 n-2 n-2 3 YM-2 = _3 S-1 M Elw +2.

Since y \in (2, \infty),

y>2 \Rightarrow y-1> 1.

le 170. Now I 0 3 70 91 thg zt ㅋ 2 Now. since 1 then JI+0. 1e Lo then 이공 뙤 +2Loo ㅋ 40thus 24 n 1 so le 1 6 (29)

Thus our assumption is right, that is, for any y in the given range, there exists x in tge the domain such that

g(x)=y.

That is g is onto.